# SpikePingpong: High-Frequency Spike Vision-based Robot Learning for Precise Striking in Table Tennis Game

### v1

@@ 5 Conclusions
In this paper, we presented SpikePingpong, a high-precision robotic table tennis system integrating spike-based vision and advanced control techniques. Our system combines SONIC, achieving millimeter-level ball-racket contact precision, and IMPACT, enabling accurate ball striking to specified target regions.. Experiments demonstrate superior performance with 91% success in the primary accuracy zone and 71% in the high-precision zone, significantly outperforming previous systems. With an inference time of just 0.407ms, SpikePingpong enables real-time decision-making while executing extended tactical sequences with a 95% completion rate. Despite these achievements, our system has limitations: it does not account for ball spin, which affects optimal interception strategies for different spin types. Additionally, performance against human players remains challenging due to the complex and unpredictable trajectories humans generate compared to our training conditions. Future work will focus on incorporating spin modeling, enhancing adaptability to diverse playing styles, and developing strategic gameplay planning for human-robot interaction. SpikePingpong advances robotic control for fast-paced tasks requiring precise spatiotemporal coordination.
@@ Appendix A Overview
Due to space limitations, we provide additional details of the proposed method in this supplementary material. In Appendix B, we present a comprehensive overview of our system architecture, including the ball launching mechanism, multi-camera vision setup, and robotic execution components. Appendix C elaborates on our ball detection and trajectory prediction methods, featuring our YOLOv4-tiny implementation, motion state estimation filtering, and hittable position prediction algorithms. In Appendix D, we describe our two specialized datasets: one for the SONIC model focusing on precise paddle-ball contact control, and another for the IMPACT model targeting effective hitting strategies. Finally, Appendix E summarizes our real-world experiments, demonstrating the system’s ability to return balls to specific target locations with high precision when tested with a ball machine, as well as video recordings of human-robot gameplay.
@@ B.2 Vision System
  • High-Speed Camera: Spike M1K40-H2-Gen3 with a resolution of 1000×1000 and a maximum frame rate of 20,000 fps. In our experiments, we set a step size of 200, saving 100 images per second, specifically designed to capture the instantaneous contact between the ball and the paddle, providing high temporal resolution image data for stroke analysis.
  • Depth Camera: Intel RealSense D455 with a resolution of 640×480 and a frame rate of 60Hz. This camera is calibrated using ArUco markers and employs YOLO [8] model for ball detection. The system transforms detected pixel coordinates into world coordinates through intrinsic and extrinsic parameters, enabling real-time tracking of the ball’s position.
@@ C.1 Ball Detection
In the context of high-speed robotic table tennis, where accurate timing and spatial awareness are critical, a high-frequency vision system is required to continuously track the ball’s position for real-time trajectory estimation and manipulation control. Thus, we chose YOLOv4-tiny [8] due to its lightweight design and computational efficiency, enabling a detection frequency of up to 150 Hz, which is crucial for the precise and timely interaction with the fast-moving ball. The initial phase of our research involves the supervised training of the detection model using publicly available datasets: Roboflow [3], TT2 [12], and Ping Pong Detection [40]. A two-phase training strategy is implemented to enhance model robustness. In the first phase, the model is trained on the complete dataset. In the second phase, samples that were misdetected during the initial training are selectively sampled and assigned higher weights for fine-tuning, with the goal of improving the detector’s accuracy on difficult instances. Our empirical validation is conducted using an Intel RealSense D455 RGB-D camera, where we employ the optimized lightweight detection model. This setup achieves a detection accuracy that exceeds 99.8%, facilitating rapid and precise 2D positional detection. The visualization results are shown in Figure 6.
@@ C.2 Motion State Estimation Filter
The system maintains a fixed-length (3 frames) history data queue for calculating initial velocity and handling data interruptions. When a data stream interruption exceeding 1 second is detected, the filter automatically resets the historical data to avoid the influence of outdated information on current estimates. This adaptive mechanism enables the system to quickly resume normal operation when the ball reappears or a new throw begins.
@@ C.3 Hittable Position Prediction
0:  $p_{0}$ (initial position), $v_{0}$ (initial velocity), $t_{current}$ (current time), $y_{plane}$ (hitting plane y-coordinate)
@@
11:  $p_{hit}\leftarrow p_{0}+v_{0}\cdot t_{hit}+\frac{1}{2}g\cdot t_{hit}^{2}$
12:  $p_{hit}.y\leftarrow y_{plane}$ Ensure exact y-coordinate
13:  $v_{hit}\leftarrow v_{0}+g\cdot t_{hit}$
14:  $t_{arrival}\leftarrow t_{current}+t_{hit}$
@@ Appendix D Dataset Description
This section details the two critical datasets developed for our table tennis robot system. These datasets support the system’s two core components: the SONIC model (Spike-Oriented Neural Improvement Calibrator) focused on precise paddle-contact control and the IMPACT model (Imitation-based Motion Planning And Control Technology) focused on effective hitting strategies. Below, we describe the collection processes, annotation methods, and how these datasets provide the foundation for the system’s performance.
D.1 Dataset for SONIC
@@ D.1.2 Data Annotation Process
[…]
@@
[…]
@@ D.1.3 Ball-Paddle Contact Detection
[…]
@@ D.2.1 Data Collection Process
After successfully training the SONIC model to ensure precise ball-paddle contact near the center of the paddle, we extended our data collection to include the hitting phase. Building upon our established interception capabilities, we implemented a structured process to collect data on effective hitting techniques. For the hitting phase data collection, we augmented our previous methodology with the following approach:
@@
    • First, intercept the ball using the SONIC model’s prediction
@@ D.2.2 Data Annotation Process
Appendix E Additional real-world experiments
We conducted extensive real-world experiments to validate the performance of our proposed system. To demonstrate the capabilities of our SpikePingpong model, we have included a demo.mp4 in the supplementary materials. This video showcases our system’s ability to return balls to different specified target locations on the opponent’s side of the table with high precision. Additionally, the video features human-robot interactive gameplay, highlighting the system’s responsiveness and adaptability in real-time competitive scenarios. These experiments confirm that our approach successfully integrates the neuromorphic vision-based ball-paddle contact detection with effective return stroke generation, resulting in a robust table tennis playing robot capable of controlled placement and responsive gameplay.

### v2

@@ 5 Conclusions
In this paper, we presented SpikePingpong, a high-precision robotic table tennis system integrating spike vision and advanced control techniques from a cognitive perspective. Our Fast-Slow system architecture emulates human visual perception processes for enhanced ball detection and trajectory prediction, while IMPACT handles strategic motion planning for accurate ball striking to specified target regions. Experiments demonstrate superior performance with 92% success in the primary accuracy zone and 70% in the high-precision zone, significantly outperforming previous systems. SpikePingpong enables real-time decision-making while executing extended tactical sequences with a 78% success rate, showcasing sustained strategic gameplay capabilities. The success of our approach stems from the Fast-Slow architecture that combines rapid physics-based prediction with neural calibration, and the IMPACT module that enables strategic ball placement through imitation learning. Beyond table tennis, the core competencies developed in this work, including high-speed object tracking, precision manipulation, and adaptive control, demonstrate broad applicability to industrial automation, medical robotics, and aerospace systems. This work advances robotic capabilities in time-critical manipulation tasks requiring precise spatiotemporal coordination.
Ethics Statement
We acknowledge and adhere to the ICLR Code of Ethics. All data collection focused solely on ball trajectories and paddle movements without storing personally identifiable information. This research contributes to recreational robotics applications and presents minimal ethical concerns when developed responsibly.
Reproducibility Statement
To ensure the reproducibility of our work, we provide comprehensive implementation details and experimental specifications throughout the paper and supplementary materials. The experimental setup, including hardware specifications, data collection protocols, and evaluation metrics, is thoroughly documented in Appendix B. The dataset collection methodology and preprocessing steps are detailed in Appendix D, enabling replication of our data generation process. Appendix E contains detailed network architectures, hyperparameters, and training procedures for all components of our system. We’ve provided the code for our method in the supplementary material. Our codebase, including model implementations, training scripts, and evaluation protocols, will be made publicly available upon publication.
Acknowledgement
This work was supported by the National Natural Science Foundation of China (62476011), and by the Beijing Natural Science Foundation (L252060).
@@ Appendix A Overview
Due to space limitations, we provide comprehensive implementation details and additional experimental validation in the appendix. Section B presents our complete system architecture, including the ball launcher, multi-camera vision setup, and robotic execution components. Section C details our real-time perception pipeline with YOLOv4-tiny ball detection, physics-based filtering, and trajectory prediction. Section D describes our two specialized datasets for the Fast-Slow System and IMPACT models. Section E provides training specifications. Section F presents extended validation, including ultra-high precision evaluation and human demonstration integration. Section G provides a comprehensive failure case analysis to identify system limitations and improvement opportunities. Section H discusses current system limitations and outlines future research directions for enhanced spin modeling, human player adaptability, and strategic gameplay planning. Section I provides a transparent disclosure of the limited use of Large Language Models in manuscript preparation. Additionally, we provide supplementary videos demonstrating the robotic system’s real-time ball-hitting performance and trajectory interception capabilities.
@@ B.2 Vision System
  • High-Frequency Camera: Spike M1K40-H2-Gen3 with a resolution of 1000×1000 and a maximum frame rate of 20,000 fps. The device is specifically configured to capture the instantaneous contact between the ball and the paddle, ensuring high temporal resolution image data for detailed stroke analysis.
  • Depth Camera: Intel RealSense D455 with a resolution of 640×480 and a frame rate of 60Hz. This camera is calibrated using ArUco markers and employs YOLO (bubbliiiing, 2020) model for ball detection. The system transforms detected pixel coordinates into world coordinates through intrinsic and extrinsic parameters, enabling real-time tracking of the ball’s position.
@@
Figure 5 illustrates the visual comparison between spike camera and conventional RGB camera captures. Due to its ultra-high frame rate, the spike camera eliminates motion blur entirely, providing crisp imagery of fast-moving objects that would appear blurred in standard cameras, which is crucial for precise ball trajectory analysis during high-speed table tennis gameplay.
Figure 5: Comparison of ball capture qualit. Conventional RGB camera with motion blur at 60 fps, Spike camera with crisp imagery at 20,000 fps, demonstrating the advantage of ultra-high frame rate for fast-moving object detection.
@@ C.1 Ball Detection
In the context of high-speed robotic table tennis, where accurate timing and spatial awareness are critical, a high-frequency vision system is required to continuously track the ball’s position for real-time trajectory estimation and manipulation control. Thus, we chose YOLOv4-tiny (bubbliiiing, 2020) due to its lightweight design and computational efficiency, enabling a detection frequency of up to 150 Hz, which is crucial for the precise and timely interaction with the fast-moving ball. The initial phase of our research involves the supervised training of the detection model using publicly available datasets: Roboflow (Alexandrova et al., 2015), TT2 (desigproject, 2022), and Ping Pong Detection (pingpong, 2024). A two-phase training strategy is implemented to enhance model robustness. In the first phase, the model is trained on the complete dataset. In the second phase, samples that were misdetected during the initial training are selectively sampled and assigned higher weights for fine-tuning, with the goal of improving the detector’s accuracy on difficult instances. Our empirical validation is conducted using an Intel RealSense D455 RGB-D camera, where we employ the optimized lightweight detection model. This setup achieves a detection accuracy that exceeds 99.8%, facilitating rapid and precise 2D positional detection. The results are shown in Figure 6.
@@ C.2 Motion State Estimation Filter
The system maintains a fixed-length (10 frames) history data queue for calculating initial velocity and handling data interruptions. When a data stream interruption exceeding 1 second is detected, the filter automatically resets the historical data to avoid the influence of outdated information on current estimates. This adaptive mechanism enables the system to quickly resume normal operation when the ball reappears or a new throw begins.
@@ C.3 Hittable Position Prediction
[…]
@@
[…]
@@ Appendix D Dataset Description
This section details the two critical datasets developed for our table tennis robot system. These datasets support the system’s two core components: the Fast-Slow System, focused on precise paddle-contact control, and the IMPACT model (Imitation-based Motion Planning And Control Technology), focused on effective hitting strategies. Below, we describe the collection processes, annotation methods, and how these datasets provide the foundation for the system’s performance.
D.1 Dataset for Fast-Slow System
@@ D.1.2 Data Annotation Process
    $\text{mm\_per\_pixel\_width}=\frac{\text{real\_width\_mm}}{\text{width\_pixels}}$ (8)
    $\text{mm\_per\_pixel\_height}=\frac{\text{real\_height\_mm}}{\text{height\_pixels}}$ (9)
@@
    $\text{mm\_per\_pixel\_avg}=\frac{\text{mm\_per\_pixel\_width}+\text{mm\_per\_pixel\_height}}{2}$ (10)
@@ D.1.3 Ball-Paddle Contact Detection
    $\text{x}_{\text{mm}}=(x_{\text{ball}}-x_{\text{paddle}})\times\text{mm\_per\_pixel}$ (12)
    $\text{y}_{\text{mm}}=(y_{\text{paddle}}-y_{\text{ball}})\times\text{mm\_per\_pixel}$ (13)
@@ D.2.1 Data Collection Process
After successfully training the Fast-Slow System to ensure precise ball-paddle contact near the center of the paddle, we extended our data collection to include the hitting phase. Building upon our established interception capabilities, we implemented a structured process to collect data on effective hitting techniques. For the hitting phase data collection, we augmented our previous methodology with the following approach:
@@
    • First, intercept the ball using the Fast-Slow System’s prediction
@@ D.2.2 Data Annotation Process
Appendix E Training Details.
Both System 2 and IMPACT modules were trained for 2000 epochs using the Adam optimizer with an initial learning rate of 1e-3 and cosine annealing schedule. System 2 used a batch size of 32 with K=10 consecutive frames as input for trajectory prediction. The IMPACT module employed a batch size of 4 to handle the complexity of return planning tasks. All positional inputs were normalized to the range [0,1], while standard scaling was applied to velocity measurements to ensure numerical stability across varying ball conditions. Training was conducted on a workstation equipped with an NVIDIA RTX 4090 GPU.
Appendix F Additional real-world experiments
F.1 Ultra-High Precision Evaluation
To further validate our system’s precision capabilities and address questions regarding our target zone selection, we conducted additional experiments using 10cm radius targets following the methodology of Büchler et al. (Buchler et al., 2022).
Experimental Setup.
We evaluated our system using 10cm radius target zones across the same four regions (A, B, C, D) used in our main experiments. This ultra-high precision threshold represents only 1.5% of the reachable table area, providing an extremely challenging benchmark for precision assessment.
Table 8: Ultra-high precision evaluation results with 10cm radius targets
Method A(%) B(%) C(%) D(%) AVG.(%)
HYSR (Buchler et al., 2022) - - - - 8
SpikePingpong 31±4 32±3 29±3 35±2 31±3
Results and Analysis.
Table 8 presents the results of our ultra-high precision evaluation. Our system achieves an average success rate of 31±3% within the 10cm target zones, substantially outperforming the 8% success rate reported by Büchler et al. (Buchler et al., 2022) using similar target specifications.
Discussion.
These results validate our choice of 20cm and 30cm target zones as meaningful precision benchmarks while demonstrating that our system maintains reasonable performance even at ultra-high precision levels. The progressive degradation from 93% (30cm) to 70% (20cm) to 31% (10cm) reflects the inherent challenges of precise ball placement in table tennis, where even human players average only 53% success in 30cm zones. Our system’s ability to achieve 31% success in 10cm targets represents a significant advancement in robotic table tennis precision control.
Appendix G Failure Case Analysis
To gain deeper insights into our system’s limitations and identify areas for improvement, we conducted a comprehensive failure analysis by categorizing all unsuccessful attempts during our precision evaluation experiments.
Table 9: Distribution of failure types in precision targeting experiments
Failure Type Percentage
Ball fails to cross net 4.6%
Correct quadrant, outside 30cm circle 79.1%
Wrong target quadrant 12.5%
Ball fails to land on table 3.8%
Results.
Table 9 presents the distribution of failure types observed in our experiments. The analysis reveals that the majority of failures (79.1%) are near-misses where the system correctly identifies the target region but falls short of the required precision threshold.
Analysis and Implications.
The failure distribution indicates robust basic control capabilities with specific areas for improvement:
  • Near-miss failures (79.1%): These cases demonstrate that our system successfully executes the fundamental task of directing the ball toward the intended table region but requires enhanced precision in fine-grained control. This suggests that improvements in trajectory optimization and control parameter tuning could yield significant performance gains.
  • Strategic errors (12.5%): Wrong quadrant targeting indicates occasional failures in high-level decision making, potentially due to perception errors or planning inconsistencies under challenging conditions.
  • Fundamental control failures (8.4%): The combined percentage of balls failing to cross the net or land on the table represents basic execution errors, suggesting room for improvement in fundamental trajectory planning and power control.
Root Cause Analysis of Near-Misses.
A deeper causal analysis reveals that the prevalence of ”near-misses” stems from two primary root causes:
  • Unmodeled Spin Dynamics: The most significant challenge is unmodeled ball spin. Extreme spin introduces complex aerodynamic effects (the Magnus effect) and alters bounce characteristics in ways our physics-based model does not capture. This can create small but critical, centimeter-level residual errors in our final trajectory prediction, often causing the ball to land just outside the target zone.
  • High Spatio-Temporal Sensitivity: The task exhibits extreme sensitivity to the precise timing and location of the ball-paddle contact. Even minor variations in the robot’s arrival time at the contact point can subtly alter the impact dynamics and the resulting ball trajectory. This high sensitivity means that even when the overall strategy is correct, slight execution imprecision can lead to a near-miss.
This analysis confirms that our system demonstrates robust basic control with the primary limitation being precision refinement rather than fundamental control failures, providing clear direction for future system enhancements.
Appendix H Limitations and Future Work
H.1 Current Limitations
Despite the achievements demonstrated in this work, our SpikePingpong system has several limitations that present opportunities for future enhancement:
Ball Spin Modeling: Our current system does not account for ball spin, which significantly affects optimal interception strategies for different spin types. The Magnus effect and varying bounce characteristics of topspin, backspin, and sidespin balls can lead to trajectory deviations that our physics-based models do not capture.
Human Player Adaptability: Performance against human players remains challenging due to the complex and unpredictable trajectories they generate compared to our controlled training conditions. Human players exhibit diverse playing styles, strategic variations, and adaptive behaviors that exceed the scope of our current training data.
H.2 Future Research Directions
Building upon the foundation established by SpikePingpong, we identify several promising directions for future research:
Spin Dynamics Integration: Incorporating comprehensive spin modeling into both the Fast-Slow system architecture and IMPACT module to handle the full spectrum of ball spin effects on trajectory prediction and strategic planning.
Adaptive Learning Framework: Enhancing the system’s adaptability to diverse playing styles through online learning mechanisms that can adjust to opponent strategies and playing patterns in real-time.
Strategic Gameplay Planning: Developing advanced strategic planning capabilities for human-robot interaction, including opponent modeling, tactical sequence planning, and adaptive game strategy formulation.
Appendix I Use of Large Language Models
Large Language Models (LLMs) were used in a limited capacity during the preparation of this manuscript. Specifically, LLMs were employed solely for:
  • Grammar correction and proofreading
  • Sentence structure improvement and clarity enhancement
  • Minor stylistic refinements to improve readability
LLMs were not involved in research ideation, experimental design, data analysis, or the generation of scientific content. All technical contributions, methodological innovations, experimental results, and scientific insights presented in this work are entirely the product of the authors’ original research.